# SpecFuse: Ensembling Large Language Models via Next-Segment Prediction

### v1

@@ Abstract
Ensembles of generative large language models (LLMs) can integrate the strengths of different LLMs to compensate for the limitations of individual models. However, recent work has focused on training an additional fusion model to combine complete responses from multiple LLMs, failing to tap into their collaborative potential to generate higher-quality responses. Moreover, as the additional fusion model is trained on a specialized dataset, these methods struggle with generalizing to open-domain queries from online users. In this paper, we propose SpecFuse, a novel ensemble framework that outputs the fused result by iteratively producing the next segment through collaboration among LLMs. This is achieved through cyclic execution of its inference and verification components. In each round, the inference component invokes each base LLM to generate candidate segments in parallel, and the verify component calls these LLMs again to predict the ranking of the segments. The top-ranked segment is then broadcast to all LLMs, encouraging them to generate higher-quality segments in the next round. This approach also allows the base LLMs to be plug-and-play, without any training or adaptation, avoiding generalization limitations. Furthermore, to conserve computational resources, we propose a model exit mechanism that dynamically excludes models exhibiting poor performance in previous rounds during each query response. In this way, it effectively reduces the number of model calls while maintaining overall performance. We conduct extensive experiments using ensembles of five LLMs with different architectures across six benchmarks, covering instruction-response, reasoning, commonsense, and instruction-following tasks. The experimental results demonstrate that SpecFuse consistently enhances performance across all benchmarks, with RougeL scores improving by $+3.1$ on the Chinese and $+3.0$ on the English human-computer interaction benchmarks. Furthermore, the model exit mechanism reduces the average models invoked per round from $5$ to $2.4$, with only a slight reduction in performance.
@@ 1 Introduction
Generative large language models (LLMs) (Brown et al., 2020; Yang et al., 2024) have been widely applied attributed to their impressive performance across various domains, providing efficient support for a broad range of user needs. These off-the-shelf generative LLMs specialize in different areas due to differences in training data and model architecture. Therefore, by combining their strengths, an ensemble of LLMs (Yang et al., 2023) can alleviate the biases and errors of individual models, delivering a better user experience. Unfortunately, vocabulary discrepancies across different LLMs limit the application of traditional logits-based fusion methods (Schapire & Freund, 2013; Sagi & Rokach, 2018) in the integration of generative LLMs.
Recent research on ensembling generative LLMs can be divided into two categories: post-hoc ensemble methods and pre-selection ensemble methods. The post-hoc ensemble method (Jiang et al., 2023b; Lv et al., 2024b) first generates complete responses for a given question by employing all base LLMs, then integrates these responses through a trained fusion model. The pre-selection ensemble method (Lu et al., 2023) pre-trains a query routing model that, for a given query, assigns it to the LLM most likely to generate a high-quality response, using only that LLM for inference.
@@
However, both methods overlook the potential for LLMs to collaboratively generate higher-quality responses through mutual inspiration during the inference process. Additionally, as these methods train an extra fusion model or routing model on specific datasets, they tend to struggle with poor generalization when faced with open-domain queries from users.
@@
In this paper, we introduce SpecFuse, a novel ensemble framework that leverages mutual inspiration between LLMs to produce high-quality next segment. Inspired by Speculative Decoding (Leviathan et al., 2023), SpecFuse achieves this by iteratively executing its two main components: Inference and Verification. In the inference component, given the preceding context, all base LLMs generate candidate fragments simultaneously, with a predefined maximum length per round. The verification component concatenates each newly generated candidate segment with the preceding context to form a batch, then feeds it into each LLM to rank the segments by calculating sequence probabilities in parallel. The top-ranked segment is then broadcast to all LLMs, inspiring them to generate higher-quality segments in the next round. In this process, there is no need to train additional fusion or routing models, which avoids generalization limitations and allows base LLMs to be effortlessly plugged in without any adaptation. Furthermore, to reduce computational costs, we propose the Model Exit mechanism, which dynamically adjusts the softmax temperature based on previous candidate rankings, modifying the distribution of cumulative model scores. Models with scores below a certain threshold will exit the response of current query, freeing up resources for other queries and reducing overall machine deployment.
@@
We select five high-performing models with 7-9 billion parameters as base LLMs and evaluate our framework across six benchmarks, covering instruction-response, reasoning, commonsense, and instruction-following tasks. Experimental results show that SpecFuse consistently enhances performance across all benchmarks, with average Rouge (n) scores improving by $+3.1$ on the English human-computer interaction benchmarks. Furthermore, the model exit mechanism reduces the average number of models invoked per round from $5$ to $2.4$, with only a slight impact on performance.
In summary, our contributions are as follows:
  • $\bullet$
@@
    We propose SpecFuse, a novel ensemble framework that generates fused results by iteratively producing the next segment through collaboration among LLMs. Our framework allows base LLMs to be effortlessly plugged in without any training or adaptation, thus avoiding generalization limitations.
  • $\bullet$
@@
  • $\bullet$
    We evaluate our framework on four tasks, including instruction-response and commonsense, and the results demonstrate that SpecFuse consistently enhances performance across six benchmarks. Additionally, the model exit mechanism reduces the average number of models invoked per round by 50%, with only minimal performance loss.
@@ 2.1 Overview
Figure 1 shows an overview of SpecFuse. Given $K$ base LLMs =$\mathcal{M}=\left\{m_{i}\right\}_{i=1}^{K}$ and an input $I$, for each round in the generation process, SpecFuse first invokes the Inference component (§ 2.2), where the base LLMs in $\mathcal{M}$ generate candidate segments in parallel. Then, it calls the Verify component (§ 2.3) to score each candidate segment, selecting the highest-scoring one as the current round output. Simultaneously, this output is concatenated with the previous input to form the new input $I$ for the next round. Finally, SpecFuse activates a Model Exit mechanism (§ 2.4), removing models with low cumulative scores from $\mathcal{M}$. The above three operations are repeated in a loop until the generated segment contains an end token.
@@ 2.2 Inference Component
Given a maximum length $L$ for the candidate segments generated at each round, the Inference component parallelly invokes each model in $\mathcal{M}$ to generate candidate segments $\left\{\mathcal{C}_{i}\right\}_{i=1}^{\left|\mathcal{M}\right|}$, extending from input $I$, and with a length not exceeding $L$, where ${\left|\mathcal{M}\right|}$ denotes the number of models in $\mathcal{M}$. The probabilities corresponding to each token in the candidate $\mathcal{C}_{i}$, are averaged to produce the self-score $\mathcal{S}{{}_{i}^{i}}$ of the model $m_{i}$ :
$\mathcal{S}{{}_{i}^{i}}=\frac{\textstyle\sum_{n=1}^{\hat{L}_{i}}(x_{i}^{n})}{\hat{L}_{i}}$ (1)
where $\hat{L}_{i}$ represents the actual length of candidate $\mathcal{C}_{i}$, as the generated candidates may be shorter than $L$ in the final round. $x_{i}^{n}$ is the probability obtained by applying Softmax normalization to the logits output by model $m_{i}$ when generating the $n$-th token in $\mathcal{C}_{i}$. Finally, each model’s generated candidate segment, together with its corresponding self-score, is input into the Verify component. We include the model’s self-score, as the do-sample [^1] https://huggingface.co/docs/transformers/generation_strategies method used by generative LLMs involves high randomness and may not yield the model’s best segment. If the model scores other candidates higher than its own output, it indicates that its generated text is of lower quality.
@@ 2.3 Verify Component
The Verify component first concatenates each candidate segment $\mathcal{C}_{i}$ with the Input $I$, forming the concatenated text $\bar{\mathcal{C}}_{i}$. These concatenated texts $\left\{\bar{\mathcal{C}}_{i}\right\}_{i=1}^{\left|\mathcal{M}\right|}$ are then grouped into a batch, with each model’s own generated candidate being removed from the batch to reduce computational load. Next, the Verify component enables all models to compute the probability of each token in the input text in parallel. Similar to obtaining $\mathcal{S}{{}_{i}^{i}}$, the probabilities of each token in segment $\mathcal{C}_{i}$ predicted by model $m_{j}$ are averaged to compute the sequence score $\mathcal{S}{{}_{i}^{j}}$, representing the evaluation of $\mathcal{C}_{i}$ by model $m_{j}$. For each candidate $C_{i}$, its self-score and the scores from other models are averaged to obtain its quality score, denoted as $\tilde{\mathcal{S}}_{i}$:
$\tilde{\mathcal{S}}_{i}=\frac{{\textstyle\sum_{j=1}^{\left|\mathcal{M}\right|}\mathcal{S}_{i}^{j}}}{\left|\mathcal{M}\right|}$ (2)
@@
In the implementation process, we use key-values cache to reduce redundant computations of previous text in both the verification and inference stages, improving the inference speed in each round. Collaborating between the inference and verification components to generate the next segment not only alleviates low-quality responses caused by a single model’s unfamiliarity with the user’s question but also reduces instability from sampling during generation. Furthermore, incorporating the best candidate segment as input for the next round can stimulate other models, and throughout the multi-round generation process, models can continuously inspire one another, ultimately leading to higher-quality responses.
@@ 2.4 Model Exit Mechanism
While model ensembles can provide users with more stable and higher-quality responses, they also come with increased computational resource demands and costs. To reduce the computational overhead without compromising performance, we propose a Model Exit mechanism. The motivation for this approach stems from our observation that, when responding to a query, some models’ output segments rarely rank first. This indicates that these models are not well-suited for responding to the given query, making further computational investment in them inefficient. We use the cumulative scores from previous rounds of each model as prior estimates of quality in subsequent rounds to determine whether a model should be exited. Since the number of rounds varies for different queries, a fixed threshold cannot be used for exit decisions. Therefore, we apply the softmax function to normalize the scores and set a temperature coefficient of $\sqrt{T}$ ($T$ being the current round). We choose $\sqrt{T}$ as the temperature coefficient because the number of output rounds rarely exceeds $100$. By using $\sqrt{T}$, we effectively limit the cumulative scores to under $10$, preventing extreme values from dominating. Additionally, we analyze the distribution of the best candidate segments. When these segments belong to only a few models, other models can be exited. By combining the temperature coefficient with the best segment distribution, the softmax scores more accurately reflect model performance, allowing underperforming models to exit promptly.
Specifically, we use $Q_{i}={\textstyle\sum_{t=1}^{T}}\tilde{\mathcal{S}}_{i}^{t}$ to represent the cumulative quality score of the candidate segments generated by model $m_{i}$ from the first step to the current $T$-th step. Next, we count the number of times each model ranked first in previous steps and weighted them based on the step intervals from the current step: within $4$ steps by $1$, $4$ to $8$ steps by $3/4$, $8$ to $12$ steps by $1/2$, and beyond $12$ steps by $1/4$, resulting in a weighted count $r_{i}$ for each model $m_{i}$. We introduced positional weights when calculating $r_{i}$ to prioritize recent steps, since models that perform well only in distant steps are less relevant to future responses. The $r_{i}$ is then normalized to create a distribution $\mathrm{P}_{r}=\left\{\tilde{r}_{i}\right\}_{i=1}^{\left|\mathcal{M}\right|}$, where $\tilde{r}_{i}$ is calculated as follows:
@@ 3 Experiments
3.1 Setups
Evaluation datasets. We evaluate all the models on six datasets that represent different core capabilities of LLMs, open-domain instruction-response (IR), commonsense, reasoning and instruction following.
  • $\bullet$
    Open-domain IR: We evaluate the model’s open-domain instruction-response capability using both English and Chinese human-computer interaction datasets. For the English dataset, we choose the Alpaca-gpt4 (Peng et al., 2023) and Dolly-15k (Conover et al., 2023) datasets for evaluation, both of which have inputs that consist of human instructions. For the Chinese dataset, we utilize the Human-Value and Ruozb datasets from the COIG-CQIA (Bai et al., 2024) benchmark for testing. A detailed description of the dataset can be found in Appendix A.
  • $\bullet$
  • $\bullet$
  • $\bullet$
@@ 3.1 Setups
Evaluation metrics. We use a variety of metrics for different tasks, following the test scripts from the Openllm leaderboard. To assess the quality of human question-answering, we apply BARTScore (Bart-S) (Yuan et al., 2021), BERTScore (Bert-S) (Zhang et al., 2019), GPT4-Rank (GPT4-R) (OpenAI et al., 2024), BLEU (Papineni et al., 2002), and ROUGE (R-n) (Lin, 2004). For multiple-choice tasks such as MMLU and ARC-C, we select the option with the highest likelihood to calculate accuracy (Acc). For the reasoning dataset GSM8K, we evaluate exact match (EM) accuracy. For IFEVAL, we rely on the evaluation files provided by the dataset creators (Zhou et al., 2023), testing under prompt-strict, instruction-strict, prompt-loose, and instruction-loose conditions. A detailed explanation of the evaluation methods is provided in Appendix B.
Baselines. Since our approach has not undergone any additional training, we selecte several types of untrained baseline models for comparison with our method: (1) Larger LLMs, including Mixtral-8x7B-v0.1 (Jiang et al., 2023a), Qwen2-72B (Yang et al., 2024), and Llama-3-70B (AI@Meta, 2024). (2) PairRank: an English reward model introduced in the LLM-Blender (Jiang et al., 2023b), which compares candidate results generated by different LLMs and selects the best candidate as the ensemble output. (3) Minimum Bayes Risk (MBR) (Freitag et al., 2023): selects the answer with the highest lexical similarity to other candidate answers. In this paper, we use the SimCSE (Gao et al., 2022) model to calculate the similarity between candidate responses. (4) Generation Fusion (GF) (Jiang et al., 2023b): uses the outputs of other models as context, passing them to a new model, which generates a response based on this context. (5) Majority Voting: each model provides a choice, and the final result is determined by the option with the most votes.
Implement details. As the methods in this paper are not trained, we only provide the parameter settings for inference. All models in this study are loaded with bfloat16 precision for inference and use the following generation parameters: $do\_sample=True$, $temperature=0.6$, and $top\_p=0.9$. All of our experiments are conducted on A100 GPUs, and set the maximum length of candidate segments to $10$.

### v2

@@ Abstract
Ensembles of generative large language models (LLMs) during inference can integrate the strengths of different LLMs to compensate for the limitations of individual models. However, recent work has focused on training an additional fusion model to combine complete responses from multiple LLMs, failing to tap into their collaborative potential to generate higher-quality responses. In this paper, we propose SpecFuse, a novel ensemble framework that outputs the fused result by iteratively producing the next segment through collaboration among LLMs during inference. This approach leverages inter-model cooperation to enhance response quality while allowing base LLMs to remain plug-and-play, requiring no training or adaptation and avoiding generalization limitations. Furthermore, to conserve computational resources, we propose a model exit mechanism that dynamically excludes models exhibiting poor performance in previous rounds during each query response. In this way, it effectively reduces the number of model calls while maintaining overall performance. We conduct extensive experiments across six benchmarks, and the experimental results demonstrate that SpecFuse consistently enhances performance across all benchmarks. Additionally, the model exit mechanism reduces the average models invoked per round from $5$ to $2.4$, with only a slight reduction in performance. We will release the code for SpecFuse.
SpecFuse: Ensembling Large Language Models via Next-Segment Prediction
Bo Lv1,2,3, Chen Tang4 , Yanan Zhang3, Xin Liu2, Yue Yu2, Ping Luo 1,2,3 1Key Lab of Intelligent Information Processing, Institute of Computing Technology, Chinese Academy of Sciences (ICT/CAS) 2Peng Cheng Laboratory 3University of Chinese Academy of Sciences 4Department of Computer Science, The University of Manchester, UK lvbo19@mails.ucas.ac.cn
@@ 1 Introduction
Figure 1: Overview of Token Logits Fusion, Complete Response Fusion, and Our Segment Fusion methods.
@@
Generative large language models (LLMs) (Brown et al., 2020; Yang et al., 2024) have been widely applied attributed to their impressive performance across various domains, providing efficient support for a broad range of user needs. These off-the-shelf generative LLMs specialize in different areas due to differences in training data and model architecture. Consequently, during the inference stage, by combining their strengths, an ensemble of LLMs (Yang et al., 2023) can alleviate the biases and errors of individual models, delivering a better user experience. Unfortunately, as shown in Figure 1 (a), vocabulary discrepancies across different LLMs and differences in tokenization methods limit the application of traditional logits-based fusion methods (Schapire and Freund, 2013; Sagi and Rokach, 2018) in the integration of generative LLMs.
To address this issue, recent work integrates the complete responses generated by multiple LLMs, as shown in Figure 1 (b). Jiang et al. (2023b) first utilize all base LLMs to generate a set of complete responses for a given query and then employ a trained ranking model to select the highest-scoring response as the ensemble output. Lv et al. (2024) train a fusion model that takes the complete responses from all models as input and integrates them to produce the final ensemble result. Similarly, Wang et al. (2024b) select a model from the base LLMs as the aggregator, then concatenate the responses from each model and re-input them into the aggregator model, which generates the fused output by referencing the other models’ outputs.
@@
However, these methods directly integrate complete responses, which not only neglect the potential for LLMs to collaboratively generate higher-quality outputs through mutual inspiration during the inference stage, but also require users to wait a long time [^1] This waiting time, known as “first token latency,” is the time users wait from inputting data into the system until the system outputs the first word. before the fused results begin to be output. Additionally, as these methods require training an extra fusion model on specific datasets or selecting a model as the aggregator, they tend to struggle with poor generalization and stability when faced with open-domain queries from users.
@@
In this paper, we introduce SpecFuse, a novel ensemble framework that leverages mutual inspiration between LLMs to produce high-quality next segment during the inference stage, as shown in Figure 1 (c). Inspired by Speculative Decoding (Leviathan et al., 2023), SpecFuse achieves this by iteratively executing its two main components: Inference and Verification. In the inference component, given the preceding context, all base LLMs generate candidate fragments simultaneously, with a predefined maximum length per round. The verification component concatenates each newly generated candidate segment with the preceding context to form a batch, then feeds it into each LLM to rank the segments by calculating sequence probabilities in parallel. The top-ranked segment is then broadcast to all LLMs, inspiring them to generate higher-quality segments in the next round. Simultaneously, it is immediately output to the user, reducing the time users wait for the base model to generate a complete response. The iterative execution of the inference and verification components eliminates the need to train a fusion model or select an aggregator, avoiding generalization and stability limitations and allowing base LLMs to be effortlessly plugged in without any adaptation.
In addition, to reduce computational costs, we propose the Model Exit mechanism, which dynamically adjusts the softmax temperature based on previous candidate rankings, modifying the distribution of cumulative model scores. Models with scores below a certain threshold will exit the response of current query, freeing up resources for other queries and reducing overall machine deployment. We select five high-performing models with 7-9 billion parameters as base LLMs and evaluate our framework across six benchmarks, covering instruction-response, reasoning, commonsense, and instruction-following tasks.
@@
Experimental results show that SpecFuse consistently enhances performance across all benchmarks, with average Rouge (n) scores improving by $+3.1$ on the English human-computer interaction benchmarks. Furthermore, the model exit mechanism reduces the average number of models invoked per round from $5$ to $2.4$, with only a slight impact on performance. In summary, our contributions are as follows:
@@
  • We propose SpecFuse, a novel ensemble framework that generates fused results by iteratively producing the next segment through collaboration among LLMs. Our framework effectively reduces first token latency and allows base LLMs to be effortlessly plugged in without any training or adaptation, thus avoiding generalization limitations.
@@
  • We evaluate our framework on four tasks, and the results demonstrate that SpecFuse consistently enhances performance across six benchmarks. Additionally, the model exit mechanism reduces the average number of models invoked per round by 50%, with only minimal performance loss.
@@ 2.1 Overview
Figure 2 shows an overview of SpecFuse. Given an input $I$ and a model list =$\mathcal{M}=\left\{m_{i}\right\}_{i=1}^{K}$ containing $K$ base LLMs, for each round in the generation process, SpecFuse first invokes the Inference component (§ 2.2), where the base LLMs in $\mathcal{M}$ generate candidate segments in parallel. Then, it calls the Verify component (§ 2.3) to score each candidate segment, selecting the highest-scoring one as the current round output. Simultaneously, this output is concatenated with the previous input to form the new input $I$ for the next round. Finally, SpecFuse activates a Model Exit mechanism (§ 2.4), removing models with low cumulative scores from $\mathcal{M}$. The above three operations are repeated in a loop until the generated segment contains an end token.
@@ 2.2 Inference Component
Given a maximum length $L$ for the candidate segments generated at each round, the Inference component parallelly invokes each model in the model list $\mathcal{M}$ to generate candidate segments $\left\{\mathcal{C}_{i}=t_{i}^{1},\cdots,t_{i}^{L}\right\}_{i=1}^{\left|\mathcal{M}\right|}$, extending from input $I$, and with a length not exceeding $L$, where ${\left|\mathcal{M}\right|}$ denotes the number of models in $\mathcal{M}$. The $j$-th token in $\mathcal{C}_{i}$ is generated by the following process:
$t_{i}^{j}=\mathcal{M}_{i}([I,t_{i}^{1}\dots t_{i}^{j-1}],\theta_{i})$ (1)
where $\mathcal{M}_{i}$ represents the $i$-th model in the model list, and $\theta_{i}$ denotes the model parameters. The probabilities corresponding to each token in the candidate $\mathcal{C}_{i}$, are averaged to produce the self-score $\mathcal{S}{{}_{i}^{i}}$ of the model $\mathcal{M}_{i}$ :
$\mathcal{S}{{}_{i}^{i}}=\frac{\textstyle\sum_{j=1}^{\hat{L}_{i}}\mathbb{P}(t_{i}^{j})}{\hat{L}_{i}}$ (2)
where $\hat{L}_{i}$ represents the actual length of candidate $\mathcal{C}_{i}$, as the generated candidates may be shorter than $L$ in the final round. $\mathbb{P}(t_{i}^{j})$ is the probability obtained by applying Softmax normalization to the logits output by model $\mathcal{M}_{i}$ when generating the $t_{i}^{j}$. Finally, each model’s generated candidate segment, together with its corresponding self-score, is input into the Verify component. We include the model’s self-score, as the do-sample [^2] https://huggingface.co/docs/transformers/generation_strategies method used by generative LLMs involves randomness and may not yield the model’s best segment. If the model scores other candidates higher than its own output, it indicates that its generated text is of lower quality.
@@ 2.3 Verify Component
The Verify component first concatenates each candidate segment $\mathcal{C}_{i}$ with the Input $I$, forming the concatenated text $\bar{\mathcal{C}}_{i}$. These concatenated texts $\left\{\bar{\mathcal{C}}_{i}\right\}_{i=1}^{\left|\mathcal{M}\right|}$ are then grouped into a batch, with each model’s own generated candidate being removed from the batch to reduce computational load. Next, the Verify component enables all models to compute the probability of each token in the input text in parallel. Similar to obtaining $\mathcal{S}{{}_{i}^{i}}$, the probabilities of each token in segment $\mathcal{C}_{i}$ predicted by model $\mathcal{M}_{n}$ are averaged to compute the sequence score $\mathcal{S}{{}_{i}^{n}}$, representing the evaluation of $\mathcal{C}_{i}$ by model $\mathcal{M}_{n}$. For each candidate $C_{i}$, its self-score and the scores from other models are averaged to obtain its quality score, denoted as $\tilde{\mathcal{S}}_{n}$:
$\tilde{\mathcal{S}}_{i}=\frac{{\textstyle\sum_{n=1}^{\left|\mathcal{M}\right|}\mathcal{S}_{i}^{n}}}{\left|\mathcal{M}\right|}$ (3)
@@
In the implementation process, we use key-values cache to reduce redundant computations of previous text in both the verification and inference stages, improving the inference speed in each round. Collaborating between the inference and verification components to generate the next segment not only alleviates low-quality responses caused by a single model’s unfamiliarity with the user’s question but also reduces first-token latency, shifting from waiting for a complete response to waiting for just one segment. Furthermore, incorporating the best candidate segment as input for the next round can stimulate other models, and throughout the multi-round generation process, models can continuously inspire one another, ultimately leading to higher-quality responses.
@@ 2.4 Model Exit Mechanism
While model ensembles can provide users with more stable and higher-quality responses, they also come with increased computational resource demands and costs. To reduce the computational overhead without compromising performance, we propose a Model Exit mechanism. The motivation for the Model Exit mechanism is detailed in the Appendix A.
Specifically, we use $Q_{i}={\textstyle\sum_{t=1}^{T}}\tilde{\mathcal{S}}_{i}^{t}$ to represent the cumulative quality score of the candidate segments generated by model $\mathcal{M}_{i}$ from the first step to the current $T$-th step. Next, we count the number of times each model ranked first in previous steps and weighted them based on the step intervals from the current step: within $4$ steps by $1$, $4$ to $8$ steps by $3/4$, $8$ to $12$ steps by $1/2$, and beyond $12$ steps by $1/4$, resulting in a weighted count $r_{i}$ for each model $\mathcal{M}_{i}$. We introduced positional weights when calculating $r_{i}$ to prioritize recent steps, since models that perform well only in distant steps are less relevant to future responses. The $r_{i}$ is then normalized to create a distribution $\mathrm{P}_{r}=\left\{\tilde{r}_{i}\right\}_{i=1}^{\left|\mathcal{M}\right|}$, where $\tilde{r}_{i}$ is calculated as follows:
@@ 3 Experiments
3.1 Experiment Setups
@@
Evaluation datasets. We evaluate all the models on six datasets that represent different core capabilities of LLMs, open-domain instruction-response (IR), commonsense, reasoning and instruction following. Please refer to Appendix B.1 for a detailed description of the dataset.
Evaluation metrics. We use a variety of metrics for different tasks, following the test scripts from the Openllm leaderboard. Please refer to Appendix B.2 for a detailed description of the evaluation methods.
Baselines. Since our approach has not undergone any additional training, we selecte several types of untrained baseline models for comparison with our method. Please refer to Appendix B.3 for a detailed description of our baselines.
Implement details. As the methods in this paper are not trained, we only provide the parameter settings for inference. Please refer to Appendix B.4 for a detailed description of our experiment implementation details.